A rocket is fired at a speed of 79.0 m/s from ground level, at an angle of 47.0 ° above the horizontal. The rocket is fired toward an 11.3-m high wall, which is located 16.0 m away. The rocket attains its launch speed in a negligibly short period of time, after which its engines shut down and the rocket coasts. By how much does the rocket clear the top of the wall?
Answer #1

Considering horizontal direction,

D = vt

16 = (79 cos47) t

Time, t = 0.30 sec

Now vertical direction,

S = ut + 0.5 at^2

S = (79 sin47) 0.30 - 0.5 x 9.8 x 0.3^2

S = 16.73 m

H = 16.73 - 11.3 = 5.43 m

Comment in case any doubt please rate my answer ...